12 forces are equal in magnitude and in equilibrium then angle between adjacent forces will be?

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
As already asked qstn. before,
​Answering again for ease:
As all forces are in equilibrium, so the net force Fnet =0.
to remain in equllibrium the 12 forces must be the sides of a regular dodecagon i.e. 12-gon (as they are equal in magnitude.
angle of a dodecagon= (n-2)*180 /n=(12-2)*180 /12
=150 degrees.
so angle between adjacent forces=150 degrees. [Ans]
 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)